Question 19

Identify the domain of the following functions.

a) 𝑓(π‘₯) = sinβˆ’1(3π‘₯ βˆ’ 1)

b) 𝑓(π‘₯) = [log(sinβˆ’1(√π‘₯2 + 3π‘₯ + 2))]

c)π‘˜(π‘₯)= 1 √(π‘₯βˆ’2)2

d) 𝑗(π‘₯)= 1 π‘₯βˆ’βˆš(π‘₯+2)

e) 𝑓(π‘₯)=ln|π‘₯+3|βˆ’5


Β 

Β Β 


1
Expert's answer
2021-09-21T05:02:10-0400

"(a) 𝑓(𝑥) = sin^{-1}(3𝑥 − 1)\\

-1<=3x-1<=1\\
\Rightarrow 0<=3x<=2\\
\Rightarrow 0<=x<=\frac{2}{3}\\
\therefore x\in [0,\frac{2}{3}]"


"(b) 𝑓(𝑥) = [log(sin^{−1}(\sqrt{𝑥^2 + 3𝑥 + 2}))]\\


\sin ^{-1} \sqrt{x^{2}+3 x+2}>0 \\
\Rightarrow \sqrt{x^{2}+3 x+2}>0 \ true \ for \ all \ \mathrm{x} . \\
Also,
 
-1 \leq \sqrt{x^{2}+3 x+2} \leq 1\\
 \Rightarrow 0<\sqrt{x^{2}+3 x+2}<=1\\
Here \ 
x^{2}+3 x+2>0 \quad \text { or } \ 
x^{2}+3 x+2 \leq 1\\
 \Rightarrow x^{2}+3 x+2>0 \quad \text { or } \ 
x^{2}+3 x+1\leq 0\\
 (x+1)(x+2)>0 \ or\ 
 
\frac{-3-\sqrt{5}}{2} \leq x \leq \frac{-3+\sqrt{5}}{2} \\
Final \ answer : {\left[\frac{-3-\sqrt{5}}{2},-2\right) \cup\left(-1, \frac{-3+\sqrt{5}}{2}\right]}"


"(c) \ 𝑘(𝑥)= \frac{1}{ \sqrt{(𝑥−2)^2}}\\
Here, \ x-2\neq0\\
\Rightarrow x\neq 2 \\
So, x\in R-{{2}}"


"(d) 𝑗(𝑥)= \frac{1}{ 𝑥−\sqrt{(𝑥+2)}}"

Here,Β "𝑥−\sqrt{(𝑥+2)}\neq0\\"

"\Rightarrow x\neq\sqrt{(𝑥+2)}\\
\Rightarrow x^2\neq x+2\\
\Rightarrow x^2-x-2\neq0\\
\Rightarrow (x-2)(x+1)\neq 0\\
\Rightarrow x\neq -1,2\\"

Also,Β 

"x+2>=0\\
x>=-2"


"\therefore x\in [-2,-1)\cup(-1,2)\cup(2,\infty)"


"(e) \ 𝑓(𝑥)=ln|𝑥+3|−5\\

Here, |𝑥+3|>0"Β which is always true.

So,Β "x\in R"